OpenStudy (anonymous):

How do you factor 2e^x - 2xe^x - 1 = 0? Can you take ln of individual terms?

OpenStudy (anonymous):

I dont think it can be factored

OpenStudy (anonymous):

@ganeshie8 can u help?

ganeshie8 (ganeshie8):

doesnt look factorable... and solving analytically looks painful hmm

OpenStudy (anonymous):

but you can solve for x here

OpenStudy (richyw):

uh...\[2e^x(1-x)=1\]

ganeshie8 (ganeshie8):

@mark_o. hw did u get that.. care to explain ?

OpenStudy (richyw):

I have no idea how you got that. I get stuck at \[e^x(x-1)=-\frac{1}{2}\]

ganeshie8 (ganeshie8):

even me.. i i take ln both sides, x-1 is going inside ln.. . i get stuck

OpenStudy (anonymous):

yup 2e^x - 2xe^x - 1 = 0 factoring 2 (1-x) -1 =0 e^x (1-x)=1/2 take ln on both sides ln [e^x (1-x)]= ln1/2 lne^x + ln(1-x)=ln1/2 x+ ln(1-x)=ln1/2

OpenStudy (anonymous):

oops sorry i thought that was lne instead of lne^x

ganeshie8 (ganeshie8):

ohhkk :)

OpenStudy (anonymous):

ok lets have fun helping some more here guys lol....